# DIRAS: Efficient LLM-Assisted Annotation of Document Relevance in Retrieval Augmented Generation

### v1

@@ 5 Recommendation for Future RAG
Given the calibrated prediction of DIRAS fine-tuned $\mathcal{M}_{o}$, an alternative way is to retrieve all documents whose relevance scores exceed a pre-defined threshold. Thus, different questions can retrieve different amounts of information depending on whether passing the relevance threshold. Advanced RAG designs can even strategically pick the calibrated threshold for different questions, for example, allowing more partial relevance for summary queries. Fig. 6 shows the F1 Scores of GPT-4 and Llama3-Tok with different relevance thresholds. Llama3-Tok achieves good F1 scores over a wide range of thresholds. Thanks to its compact size (8B), it can be efficiently deployed as a reranker in RAG systems.
@@
6 Related Literature
IR plays an important role in RAG but also becomes a performance bottleneck (Gao et al., 2024). Low precision in IR may cause LLMs to hallucinate or pick up irrelevant information (Cuconasu et al., 2024; Schimanski et al., 2024a). Low recall may leave out critical information for analysis (Ni et al., 2023). Prior work proposes various datasets and algorithms to benchmark RAG performance (Saad-Falcon et al., 2023; Chen et al., 2023a; Niu et al., 2024), but most of them focus on the final generation quality. No previous work studies how to efficiently generate domain-specific IR benchmarks for RAG. Sun et al. (2023b, a); Pradeep et al. (2023); Qin et al. (2024) find that SOTA generic LLMs are good rerankers and such ability can be distilled to open-sourced LLMs. However, they all focus on pairwise or listwise ranking methods, while our work shows that the pointwise method (1) better fulfills the need of annotating domain-specific IR; and (2) works better with proper calibration method (Tian et al., 2023). Our work follows the stream of research annotating document relevance with LLMs (Thomas et al., 2024), taking one step further by (1) showing how to annotate with small LLMs; (2) predicting relevance scores to address partial relevance; and (3) use them to benefit RAG development.
@@ 7 Conclusion
In this work, we introduce the DIRAS pipeline to fine-tune open-source LLMs to calibrated annotators. We prove the effectiveness of the approach on two dataset sets. The DIRAS approach has two significant advantages: (1) it is case-specialised allowing the incorporation of domain-specific knowledge into definitions, and (2) it helps to efficiently label a huge amount of documents with calibrated relevance scores.
@@
As with every work, this has limitations. Our first limitation stems from the usage of two datasets focusing on a scenario of RAG report analyses. Given our expertise, this allows us to extend the deepness of our investigations: annotating domain-specific benchmarks and conducting error analyses. However, this limits the wideness of our research. While we argue that the results are representative of other knowledge-intensive RAG scenarios, it remains an open question for future work to generalize the DIRAS pipeline.
@@
Our third limitation, and also a viable option to address multimodality, lies in the recent introduction of long-context LLMs. These may make the role of information retrieval in RAG less crucial as entire documents can be used to answer a question. At the same time, we observe that long-context models are good in needle-in-a-haystack problems but not as good when multiplied needles exist (Team et al., 2024a). Thus, even for long-context LLMs, an efficient system like DIRAS could enable improving algorithms for finding and using multiple relevant pieces of information or help improve the model’s ability to do so.
@@ Appendix A Exemplifying Partial Relevance
Furthermore, conclude as to whether this is relevant to answer the following question and definition:
@@
The question is clearly looking for processes to identify and assess risks associated with climate change. Example 1. states that "climate risk assessment tools" are relevant. The paragraph states that the Integrated Management System serves to identify risks including environmental risks. In sustainability matters, climate change and environmental topics often fall under the same umbrella. Thus, yes, the paragraph is relevant for the question addressing a certified process to manage climate risks. However, also contrary arguments can be considered. We don’t exactly know whether environmental and climate topics are viewed interchangeably. An expert may know clear differentiating factors between environmental and climate matters (e.g., not all environmental problems like water pollution affect the climate). Furthermore, the environmental management system is rather a minor note in this paragraph. Additionally, it seems that, although it is a general risk management system, the "Quality, Environmental and Occupational Health & Safety (QEHS) Management System" is rather used to achieve worldwide objectives for the company. Would you deem this relevant if it was the only information obtained for a company? And what if there are fifteen more documents that are clearly relevant? How would it be labeled then? It is possible to go to lengths and depending on which expert level or context a labeler holds. In a binary relevant/irrelevant setting, both labels would be partially wrong. The reason lies in the fact that when asking whether this document is relevant to the question, the answer is "partially right".
Appendix B Creation of Question Relevance Definition
Fig. 7 shows the prompt template for the creation of the question relevance definition. We ask the model to produce a short definition on which the model should rely. Additionally, we ask the model to produce a list of examples. This structure should align with the manner an expert implicity or explicitly approaches the annotation task of labeling relevance. A definition alone would have the shortcoming that it only incorporated generic know-how. Complementing it with examples gives the expert the flexibility to extend the meaning of the terms in exemplified form. For a demonstration of the output, see App. J.
An analyst posts a <question> about a sustainability report. Your task is to explain the <question> in the context of sustainability reporting. Please first explain the meaning of the <question>, i.e., the meaning of the question itself and the concepts mentioned. And then give a list of examples, showing what information from the sustainability report the analyst is looking for by posting this <question>.
@@ Appendix B Creation of Question Relevance Definition
Figure 7: RAG Prompt Template enforcing structured output.
@@ Appendix C Metrics Computation Details
In this project, we use Scikit-Learn (version 1.2.2) to compute AUROC, average precision scores, Brier scores, and F1 scores. We employ rank_eval (version 0.1.3) to compute nDCG and MAP scores, and Scipy.stats to compute Kendall’s $\tau$. For nDCG, relevant scores 0.5 and 1 are assigned to partially relevant and relevant documents correspondingly.
Appendix D PDF Parsing and Document Length
We use IBM deepsearch parser (Team, 2022) to parse corporate reports into chunks. For chunks shorther than 120 tokens, we concatenate them with adjacent chunks to form chunks longer than 120. Figure Fig. 8 shows the formatted chunks length distribution.
Figure 8: Distribution of chunk length after being extracted from sustainability reports and concatenation.
@@ Appendix E Expert Annotation Process
To obtain an expert-annotated test set, we annotate the test data that contains equally 330 samples of each top-5 relevant and non-to-5 relevant documents to the 11 questions/queries, thus obtaining 660 (query, document) pairs including the corresponding expert definitions. As described in App. D, the data is obtained from real sustainability reports and split into chunks of around 150 words with the IBM deepsearch parser (Team, 2022). Table 6 shows an overview of statistical properties of the number of words in test set data.
@@
Table 6: Statistical properties of the number of words in ChatReport test set data.
@@
  • If a paragraph clearly fall into the definition of relevance, i.e., explicitly mentioned by the question explanation or examples, please annotate relevant.
@@
Additionally, one PhD student focusing on climate change and sustainability research serves as a subject-matter meta annotator to resolve conflicts or investigate cases where both labelers arise at the label partially.
@@
Table 7: Label distribution in the ChatReport test dataset.
@@ Appendix G DIRAS Prompt Template $\mathcal{P}$
You are a helpful assistant who assists human analysts in identifying useful information within sustainability reports for their analysis.
@@ Appendix H Alternative Prompts
Appendix I ClimRetrieve Dataset Overview
The ClimRetrieve dataset simulates the typical tasks of a sustainability analyst. The annotators examine 30 sustainability reports. We use the report-level dataset of this paper which contains 43K (query, document) pairs labelled by relevance. This dataset is very long because every report is repeated per question to have unique (query, document) pairs. We also use the much shorter, only relevant (query, document) pairs containing 595 unique samples. Since the annotators only search for relevant information, these are the gold labels. There is no active labeling of the irrelevant (query, document) pairs. Table 8 offers a comparison of the statistical properties of the word count of the documents in ClimRetrieve. They are slightly longer than the ChatReport data (compare Table 6). Table 9 shows the label distribution in the relevant-only dataset. It becomes apparent that most of the (query, document) pairs are very relevant (label=3). This aligns with the report analyst setting where information is searched for until the question can effectively be answered.
Number of words per document
Dataset Dataset Size Mean Std Min 25% 50% 75% Max
Report-Level 43.445 172 60.0 1 135 186 220 499
Only Relevant 595 199 43.8 38 177 213 228 281
Table 8: Statistical properties of the number of words in ClimRetrieve data.
Label Occurance
1 100
2 167
3 328
Table 9: Distribution of relevance labels over the relevant (query, document) pairs in ClimRetrieve.
Appendix J Creation of the Expert-Informed Relevance Definition
Fig. 14 shows the prompt for the creation process of the expert-informed relevance definitions. Following the procedure in Schimanski et al. (2024b), we make use of the text parts labeled as relevant. There exists a relevance score from 1-3 where 1 signals the least and 3 is most relevant. Similar to the base setup for the experiments in Schimanski et al. (2024b), we use the text samples with a score of 2 or higher to create the expert-informed relevance definition. We include every relevant text part as an example. Thus, the prompt includes a large set of examples per question that can be synthesized in the relevance definition. While this procedure creates an expert-informed definition, it also represents a data leakage. In a sense, the definition will retrofit with the documents it searches for. While this is a limitation, we argue that this is also the step an expert human would take when annotating. She will have fixed concepts in her head, maybe even inspired by prior search processes. Therefore, we argue that this data leakage experiment is still adequate to investigate whether an expert-informed definition helps align the search process.
Plugging the examples into the prompt results in a set of expert-informed relevance relevance definitions. When comparing these relevance definitions to the generic ones, it becomes apparent that GPT-4 already incorporated the majority of the concepts that the experts were looking for. Therefore, the adjustment of the relevance definition is visible but rather subtle. One example is displayed in App. J. While the meaning of the question remains rather static, there are nuanced differences in the examples that guide the relevance labeling.
An analyst posts a <question> about a sustainability report. Your task is to explain the <question> in the context of sustainability reporting. Please first explain the meaning of the <question>, i.e., meaning of the question itself and the concepts mentioned. And then give a list of examples, showing what information from the sustainability report the analyst is looking for by posting this <question>.
@@ Appendix I ClimRetrieve Dataset Overview
Figure 14: RAG Prompt Template enforcing structured output with the inclusion of examples.
@@
Does the company provide definitions for climate change adaptation? "Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly outlined what it understands by the term ""climate change adaptation"" within its sustainability reporting. This involves understanding how the company perceives the process of adjusting to actual or expected climate change and its effects, as well as how it plans to manage or mitigate harm or exploit beneficial opportunities. Examples of information that the question is looking for: 1. A clear definition of climate change adaptation as understood and applied by the company in its operations and planning. 2. Descriptions of the specific impacts of climate change that the company is adapting to, such as increased frequency of extreme weather events, rising sea levels, or changing agricultural conditions. 3. Examples of measures the company is taking to adapt to these impacts, such as investing in flood defenses, altering supply chain logistics, or developing drought-resistant crops. 4. Information on how the companys definition of climate change adaptation aligns with or differs from standard definitions provided by authoritative bodies like the Intergovernmental Panel on Climate Change (IPCC) or the United Nations Framework Convention on Climate Change (UNFCCC). 5. Details on any collaborative efforts with governments, NGOs, or other corporations to address climate change adaptation collectively. 6. Insights into how the company measures or evaluates the effectiveness of its adaptation strategies." "Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly defined what it means by climate change adaptation within its sustainability reporting. This involves understanding how the company interprets climate change adaptation, including any specific strategies, initiatives, or policies it has in place to adjust to current or expected future climate change impacts. Examples of information that the question is looking for: 1. Descriptions of how the company defines climate change adaptation in the context of its operations and strategic planning. 2. Examples of specific adaptation measures the company has implemented or plans to implement, such as enhancing infrastructure resilience, diversifying water sources, or adjusting agricultural practices. 3. Information on how the companys definition of climate change adaptation aligns with or diverges from standard definitions provided by environmental organizations or regulatory bodies. 4. Details on how the company assesses and integrates climate change risks and opportunities into its investment decision-making processes, focusing on adaptation. 5. Statements on the companys involvement in partnerships or alliances aimed at promoting climate change adaptation and resilience, indicating a collaborative approach to defining and addressing adaptation needs."
@@
Appendix K MAP and nDCG Scores for Different Relevance Definitions
In the expert-informed definition experiment, we compare two settings. First, we compare the predictions on the 595 relevant-only (query, document) pairs. This is a replication of the setting in § 4.1. Since we don’t have non-relevant samples, we can only compare the nDCG. Table 11 shows the results. It becomes apparent that only for the general nDCG score, the expert-informed query rates better. For the nDCG@5, and nDCG@10, the best-performing model remains with the generic prompt. The picture turns again when widening to nDCG@15. This could be a result of the definition creation. We use examples of relevance labels 2 and 3 to create the expert-informed definition. Thus, we implicitly equalize relevance 2 and 3 in importance. This means we are likely less effective in differentiating between 2 and 3. This could explain the lower results for at lower k’s where differentiating between 2 and 3 is important v.s. the overall nDCG where differentiating between 1 and 2/3 plays a more important role.
@@
Appendix L Embedding Fine-Tuning
@@
We follow the official fine-tuning example [^9] https://github.com/FlagOpen/FlagEmbedding/tree/master/examples/finetune of (Chen et al., 2024) to fine-tune the embedding models. The models are fine-tuned on all annotated (query, document) pairs in ChatReport test set for 10 epochs, with a batch size of 4. Other hyperparameters are the same as the official example.
@@
Appendix M Hand-Checking of Model Results on ClimRetrieve
The ClimRetrieve dataset presents the opportunity to investigate the output of our best-performing fine-tuned model Llama3 without CoT from different angles. In this experiment, we hand-annotate the output of our model in different categories and compare it to the ClimRetrieve annotations. As outlined in § 4, one issue of our approach is that we can only approximate the mental model of the annotator to create the relevance definitions. For the analysis, we consider the predictions of the fine-tuned Llama-3 model with the expert-informed relevance definitions created in App. J as they likely present the closest approximation to the mental model of the labelers. Furthermore, the ClimRetrieve dataset only has golden labels for the (partially) relevant documents. This has implications for our results. In this setup, we view our model as a calibrated annotator acting according to the given relevance definition – i.e., our model represents a fictional golden truth. This allows us to perform a qualitative edge case analysis on the ClimRetrieve data in different categories.
The first category is the true positive classifications (model says relevant, ClimRetrieve-human says relevant). Since they are golden annotations by the ClimRetrieve annotators and align with our model, they are not checked as they likely are error-free. A more nuanced view has to be adopted for the three categories false negatives (model says relevant, ClimRetrieve-human says not relevant), true negatives (model says not relevant, human says not ClimRetrieve-relevant), and false positives (model says not relevant, human says ClimRetrieve-relevant). To obtain a qualitative understanding, we investigate the appearances (see Table 13) and confidences (see Table 14) of the categories. For the confidences, we use the empirically best-performing model in information retrieval Llama3-Tok (see Table 2). To perform a qualitative investigation, we sample (document, query) pairs from each category and reannotate them as being "relevant", "partially relevant" or "not relevant". These reannotations do not serve to obtain a holistic view (as already done in § 3) but rather gain insights into the special cases of the model’s predictions.
First, we investigate the false negatives (model says relevant, human says not relevant). Investigating these cases is of particular interest since we know that humans have a selection bias (Thakur et al., 2021). In the analyst scenario setup of ClimRetrieve, this circumstance is aggravated by the fact that analysts may only search for information until they deem that they can answer the question at hand. Looking at Table 13, it becomes apparent that there is a very large number of false negatives. However, Table 14 shows that the confidence in the true negatives is much higher than in the false negatives. This indicates that the false negatives contain a much larger spectrum of partial relevance. This can be an initial explanation for the inflated number of false negatives.
To perform a qualitative investigation, we sample (document, query) pairs from the false negatives and perform a reannotation with a single annotator. We have 16 unique questions/queries in the ClimRetrieve dataset and sample 6 (document, query) pairs per question/query. Furthermore, we want to investigate which role model confidence plays. Thus, the (document, query) pairs are sampled to contain two samples with a Llama3-Ask confidence above 0.9, two between 0.9 and 0.7, and two below 0.7. We choose Llama3-Ask confidence because it is easier to form thresholds and approximates the confidence well enough.
Since we know that our model and the ClimRetrieve annotation inherently disagree, we focus on our reannotation. Fig. 16 shows the resulting percentage of agreement between the model and our reannotation for increasing the confidence threshold. It becomes apparent that irrespective of the threshold, the agreement between our reannotator and the model is very high. Furthermore, increasing the threshold correlates with a high model-reannotation agreement. This indicates that the model’s assignment of the "relevant" label indeed works. At the same time, it bears the question of whether the initial ClimRetrieve-human annotation is invalid. Why does it underestimate the true relevant documents by such a large magnitude? There are three avenues to explain these results. First, as already mentioned, humans may miss out on sources, suffer from selection bias, not taking duplicate information into account, and may stop annotating once they reach a sufficient level of information for a question. While these things are inherent in the analyst setting of ClimRetrieve and out of our control, we can observe that ClimRetrieve clearly misses out on some relevant information (see for instance Fig. 15). However and second, one further major reason for the inflated number of false negatives may be the change of definition we make by creating it with GPT-4. The original questions all concern the specific topic of climate change adaptation. This is a very narrow, specialized case of the general climate change domain. In our definitions, this generally looks different. Consider for instance the following example:
@@
Meaning of the question: The question "Do the environmental/sustainability targets set by the company reference external climate change adaptation goals/targets?" is asking for information about whether the companys stated goals or objectives for environmental sustainability or climate change mitigation are aligned with, or make reference to, established external goals or targets. These external references could include international agreements, national policies, or standards set by recognized organizations focused on climate change and sustainability.

### v2

@@ 5 Recommendation for Future RAG
Given the calibrated prediction of DIRAS $\mathcal{M}_{s}$, an alternative way is to retrieve all documents whose relevance scores exceed a pre-defined threshold. Thus, different questions can retrieve different amounts of information depending on whether passing the relevance threshold. Advanced RAG designs can even strategically pick the calibrated threshold for different questions, for example, allowing more partial relevance for summary queries. Fig. 6 shows the F1 Scores of GPT-4 and Llama3-Tok with different relevance thresholds. Llama3-Tok achieves good F1 scores over a wide range of thresholds. Thanks to its compact size (8B), it can be efficiently deployed as a reranker in RAG systems.
@@
6 Background and Related Work
IR plays an important role in RAG but also becomes a performance bottleneck (Gao et al., 2024). Low precision in IR may cause LLMs to hallucinate or pick up irrelevant information (Cuconasu et al., 2024; Schimanski et al., 2024a). Low recall may leave out critical information for analysis (Ni et al., 2023). Domain-specific knowledge is also important for retrieval performance (Tang and Yang, 2024).
Prior work on IR in RAG has already explored the idea of using LLMs to judge relevance. The closest to our work are RAGAs and ARES. Es et al. (2024) develop RAGAs to evaluate the relevance of each sentence in a paragraph with a closed-source LLM and create an aggregated score by dividing the number of relevant sentences over all sentences. Saad-Falcon et al. (2023b) aim to bring a relevance judge to a target domain through few-shot in-context learning (ICL): they first generate synthetic questions to given target-domain passages, and then fine-tune small classifiers for relevance judgment. However, many real-world challenges remain unaddressed. Specifically, IR recall, partial relevance, and domain-specific relevance definitions are neglected. ARES appears to have target-domain IR evaluation, but the synthetic data approach focuses on less integrative queries: each question is generated given a single passage and hard negatives are passages sampled from the same document, which could be (partially) relevant for integrative queries asking for broader information. Furthermore, we also find that few-shot ICL fails to teach domain-specific relevance to LLMs (§ 3.2).
Besides, Sun et al. (2023b, a); Pradeep et al. (2023); Qin et al. (2024) find that SOTA generic LLMs are good rerankers and such ability can be distilled to open-sourced LLMs. These studies all focus on pairwise or listwise ranking methods, and discusses that pointwise methods may not work due to bad calibration (Qin et al., 2024). However, when it comes to relevance annotation instead of reranking, pointwise relevance prediction is more suitable because it: (1) analyzes and annotates (query, document) pairs one-by-one, thus is more efficient and may better consider relevance definitions; and (2) can annotate both document rank and binary relevance labels (relevant or irrelevant) which are important for RAG in order to decide what information should be passed to the generator. We also show pointwise annotation works better with proper calibration method (Tian et al., 2023).
@@ 7 Conclusion
In this work, we introduce the DIRAS pipeline to fine-tune open-source LLMs to calibrated annotators. The DIRAS approach has two significant advantages: (1) it is case-specialised allowing the incorporation of domain-specific knowledge into definitions, and (2) it helps to efficiently label a huge amount of documents with calibrated relevance scores.
@@
As with every work, this has limitations. First, our results show that DIRAS fine-tuning grants small student LLMs GPT-4-level performance on specific domains, but GPT-4 is not guaranteed to be perfect in all domains and cases. In certain niche domains, it might be necessary to augment GPT-4 with domain knowledge or agentic designs to achieve human-level performance in relevance annotation, and then create reliable training data for DIRAS. Although performance not always guaranteed, LLM annotation for document relevance is still necessary due to the sheer volume of (query, document) pairs and selection bias of human annotation.
@@
Our third limitation, and also a viable option to address multi-modality, lies in the recent introduction of long-context LLMs. These may make the role of information retrieval in RAG less crucial as entire documents can be used to answer a question. At the same time, we observe that long-context models are good in needle-in-a-haystack problems but not as good when multiplied needles exist (Team et al., 2024a). Thus, even for long-context LLMs, an efficient system like DIRAS could enable improving algorithms for finding and using multiple relevant pieces of information or help improve the model’s ability to do so.
@@ Appendix A Exemplifying Partial Relevance
Furthermore, conclude as to whether this is relevant to answer the following query and definition:
@@
The query clearly looks for processes to identify and assess risks associated with climate change. Example 1. states that "climate risk assessment tools" are relevant. The paragraph states that the Integrated Management System serves to identify risks including environmental risks. In sustainability matters, climate change and environmental topics often fall under the same umbrella. Thus, yes, the paragraph is relevant for the question addressing a certified process to manage climate risks. However, also contrary arguments can be considered. We don’t exactly know whether environmental and climate topics are viewed interchangeably. An expert may know clear differentiating factors between environmental and climate matters (e.g., not all environmental problems like water pollution affect the climate). Furthermore, the environmental management system is rather a minor note in this paragraph. Additionally, it seems that, although it is a general risk management system, the "Quality, Environmental and Occupational Health & Safety (QEHS) Management System" is rather used to achieve worldwide objectives for the company. Would you deem this relevant if it was the only information obtained for a company? And what if there are fifteen more documents that are clearly relevant? How would it be labeled then? It is possible to go to lengths and depending on which expert level or context a labeler holds. In a binary relevant/irrelevant setting, both labels would be partially wrong. The reason lies in the fact that when asking whether this document is relevant to the question, the answer is "partially right".
Appendix B Creation of the Relevance Definition
Fig. 7 shows the prompt template for the creation of the query relevance definition. We ask the model to produce a short definition on which the model should rely. Additionally, we ask the model to produce a list of examples. This structure should align with the manner an expert implicity or explicitly approaches the annotation task of labeling relevance. A definition alone would have the shortcoming that it only incorporated generic know-how. Complementing it with examples gives the expert the flexibility to extend the meaning of the terms in exemplified form. For a demonstration of the output, see App. I.
An analyst posts a <question> about a climate report. Your task is to explain the <question> in the context of climate reporting. Please first explain the meaning of the <question>, i.e., the meaning of the question itself and the concepts mentioned. And then give a list of examples, showing what information from the climate report the analyst is looking for by posting this <question>.
@@ Appendix B Creation of the Relevance Definition
Figure 7: Prompt for generating a query relevance definition.
@@ Appendix C Metrics Computation Details
In this project, we use Scikit-Learn (version 1.2.2) to compute AUROC, average precision scores, Brier scores, and F1 scores. We employ rank_eval (version 0.1.3) to compute nDCG and MAP scores, and Scipy.stats to compute Kendall’s $\tau$. For nDCG, relevant scores 0.5 and 1 are assigned to partially relevant and relevant documents correspondingly. When averaging Calibration metrics, we average AUROC with $1-$ ECE and $1-$ Brier Score to keep the trend consistent.
Appendix D ChatReportRetrieve Data Preprocessing
Data Sampling: We sample 31 questions and 80 reports from this application. Importantly, ChatReportRetrieve differs from general QA dataset and focuses on integrative queries which usually ask for broad relevant information. Therefore, with controlled annotation budget (i.e., number of (query, document) pairs), we are prone to have fewer (but representative) queries and more documents for each query. The queries are strategically sampled to ensure representativeness and diversity. Specifically, 11 queries are the core queries used in ChatReport, which cover essential topics of sustainability disclosure. 20 questions are selected from users’ customized questions posed to the ChatReport tool. Climate reports are sampled randomly from openly accessible user submissions [^8] See https://github.com/EdisonNi-hku/chatreport. Finally, we prompt GPT-4 to draft relevance definitions for all queries (see App. B).
PDF Parsing: We use IBM deepsearch parser (Team, 2022) to parse corporate reports into chunks. For chunks shorther than 120 tokens, we concatenate them with adjacent chunks to form chunks longer than 120. Figure Fig. 8 shows the formatted chunks length distribution.
Figure 8: Distribution of chunk length after being extracted from climate reports and concatenation.
Train-Test Split: We split the questions into 11 for testing and 20 for training. Similarly, we split the reports into 30 for testing and 50 for training. This ensures the evaluation on unseen queries and reports. For each query, we randomly sample 60 documents – 30 each from in the top-5 and outside the top-5 (using OpenAI text-embedding-3-small as the dense retriever). Ultimately, (query, document) pairs in training split are used to create training data with relevance label and confidence score predictions (details in § 3.2). Data points in the test split are passed to human annotation. We use 31 climate-related queries for data sampling among 80 climate reports. For the separation of the train dataset and test dataset, these queries are classified into 3 categories: vague, specific, and TCFD. Within each category, queries are split randomly for training and testing, resulting in 20 queries for the train set and 11 queries for the test set. Meanwhile, the 80 climate reports are randomly split into 50 for train and 30 for test.
@@ Appendix E Expert Annotation Process
As described in App. D, the data is obtained from real climate reports and split into chunks of around 150 words with the IBM deepsearch parser (Team, 2022). Table 9 shows an overview of statistical properties of the number of words in test set data.
@@
Table 9: Statistical properties of the number of words in ChatReportRetrieve test set data.
@@
  • If a paragraph clearly falls into the definition of relevance, i.e., explicitly mentioned by the question explanation or examples, please annotate relevant.
@@
Additionally, one PhD student focusing on climate change and sustainability research serves as a subject-matter meta-annotator to resolve conflicts or investigate cases where both labelers arise at the label partially.
@@
Table 10: Label distribution in the ChatReportRetrieve testset.
@@ Appendix G DIRAS Prompt Template $\mathcal{P}$
You are a helpful assistant who assists human analysts in identifying useful information within climate reports for their analysis.
@@ Appendix H Alternative Prompts
Appendix I Creation of the Improved Relevance Definitions
Fig. 14 shows the prompt for the creation process of the improved relevance definitions. Following the procedure in Schimanski et al. (2024b), we make use of the text parts labeled as relevant. There exists a relevance score from 1-3 where 1 signals the least and 3 is most relevant. Similar to the base setup for the experiments in Schimanski et al. (2024b), we use the text samples with a score of 2 or higher to create the improved relevance definition. We include relevant text samples in the prompt for creating the relevance definitions to obtain improved definitions. The logic behind this definition creation is that we assume we know the mental model of ClimRetrieve human analysts, and thus know what information is relevant before annotation. This is common in corporate report analysis where experts will have fixed concepts in their heads, maybe even inspired by prior search processes.
Plugging the examples into the prompt results in a set of improved relevance definitions. When comparing these relevance definitions to the generic ones, it becomes apparent that GPT-4 already incorporated the majority of the concepts that the experts were looking for. Therefore, the adjustment of the relevance definition is visible but rather subtle. One example is displayed in App. I. While the meaning of the question remains rather static, there are nuanced differences in the examples that guide the relevance labeling.
An analyst posts a <question> about a climate report. Your task is to explain the <question> in the context of climate reporting. Please first explain the meaning of the <question>, i.e., meaning of the question itself and the concepts mentioned. And then give a list of examples, showing what information from the climate report the analyst is looking for by posting this <question>.
@@ Appendix I Creation of the Improved Relevance Definitions
Figure 14: Prompt Template enforcing structured output with the inclusion of examples.
@@
Does the company provide definitions for climate change adaptation? "Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly outlined what it understands by the term ""climate change adaptation"" within its climate reporting. This involves understanding how the company perceives the process of adjusting to actual or expected climate change and its effects, as well as how it plans to manage or mitigate harm or exploit beneficial opportunities. Examples of information that the question is looking for: 1. A clear definition of climate change adaptation as understood and applied by the company in its operations and planning. 2. Descriptions of the specific impacts of climate change that the company is adapting to, such as increased frequency of extreme weather events, rising sea levels, or changing agricultural conditions. 3. Examples of measures the company is taking to adapt to these impacts, such as investing in flood defenses, altering supply chain logistics, or developing drought-resistant crops. 4. Information on how the companys definition of climate change adaptation aligns with or differs from standard definitions provided by authoritative bodies like the Intergovernmental Panel on Climate Change (IPCC) or the United Nations Framework Convention on Climate Change (UNFCCC). 5. Details on any collaborative efforts with governments, NGOs, or other corporations to address climate change adaptation collectively. 6. Insights into how the company measures or evaluates the effectiveness of its adaptation strategies." "Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly defined what it means by climate change adaptation within its climate reporting. This involves understanding how the company interprets climate change adaptation, including any specific strategies, initiatives, or policies it has in place to adjust to current or expected future climate change impacts. Examples of information that the question is looking for: 1. Descriptions of how the company defines climate change adaptation in the context of its operations and strategic planning. 2. Examples of specific adaptation measures the company has implemented or plans to implement, such as enhancing infrastructure resilience, diversifying water sources, or adjusting agricultural practices. 3. Information on how the companys definition of climate change adaptation aligns with or diverges from standard definitions provided by environmental organizations or regulatory bodies. 4. Details on how the company assesses and integrates climate change risks and opportunities into its investment decision-making processes, focusing on adaptation. 5. Statements on the companys involvement in partnerships or alliances aimed at promoting climate change adaptation and resilience, indicating a collaborative approach to defining and addressing adaptation needs."
@@
Appendix J MAP and nDCG Scores for Different Relevance Definitions
In the improved definition experiment, we compare two settings. First, we compare the predictions on the 595 relevant-only (query, document) pairs. This is a replication of the setting in LABEL:sec:climre_veri. Since we do not have non-relevant samples, we can only compare the nDCG. Table 12 shows the results. It becomes apparent that only for the general nDCG score, the improved query performs better. For the nDCG@5, and nDCG@10, the best-performing model remains with the generic prompt. The picture turns again when widening to nDCG@15. This could be a result of the definition creation. We use examples of relevance labels 2 and 3 to create the improved definition with GPT-4. Thus, we implicitly equalize relevance 2 and 3 in importance. This means we are likely less effective in differentiating between 2 and 3. This could explain the lower results at lower k’s where differentiating between 2 and 3 is important v.s. the overall nDCG where differentiating between 1 and 2/3 plays a more important role.
@@
Appendix K Embedding Fine-Tuning
@@
We follow the official fine-tuning example [^10] https://github.com/FlagOpen/FlagEmbedding/tree/master/examples/finetune of (Chen et al., 2024) to fine-tune the embedding models. The models are fine-tuned on all annotated (query, document) pairs in ChatReportRetrieve test set for 10 epochs, with a batch size of 4. Other hyperparameters are the same as the official example.
@@
Appendix L Hand-Checking DIRAS Model Disagreements with ClimRetrieve
We want to check samples where our student model Llama3-Tok’s prediction differs from the relevance label in ClimRetrieve. We sample equally from those samples where Llama3-Tok indicated relevance and ClimRetrieve does not and vice versa. While our focus lies on those documents that were not annotated in ClimRetrieve (i.e. irrelevant ones), the dataset also allows us to investigate how our model performs in the edge case of when a ClimRetrieve annotator deems relevance and Llama3-Tok does not.
On the samples that were labeled as irrelevant in ClimRetrieve, we find that our student model Llama3-Tok model is effective in mitigating annotation selection bias and therefore incorporates the perspective of IR recall (see Table 5).
However, it is interesting that our student model Llama3-Tok successfully overrules human decisions. We attribute this to the fact that our created relevance definitions might differ from the mental model of the human annotator in ClimeRetrieve. Thus, humans in ClimRetrieve might have been consistent with their own mental model. For us, however, it is more important and reaffirming to observe that Llama3-Tok is consistent with its own, explicit relevance definitions.
We can view the (query, definition, document) pair in Fig. 15 as an example. When analyzing the query "Do the environmental/sustainability targets set by the company reference external climate change adaptation goals/targets?", the ClimRetrieve labeler interpreted the question broader, i.e., deeming this as relevant: "As a global technology leader, we are also committed to helping build the enabling societal conditions that will support a net zero economy.". However, for our student model, it is in line with the definition to assign a "not relevant" label. There is no explicit standard mentioned in the document.
QUERY: "Do the environmental/sustainability targets set by the company reference external climate change adaptation goals/targets?"
@@
QUERY DEFINITION: Meaning of the question: The question "Do the environmental/sustainability targets set by the company reference external climate change adaptation goals/targets?" is asking for information about whether the companys stated goals or objectives for environmental sustainability or climate change mitigation are aligned with, or make reference to, established external goals or targets. These external references could include international agreements, national policies, or standards set by recognized organizations focused on climate change and sustainability.